OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3 ln x - 2 ln c

OpenStudy (anonymous):

@wolf1728 @whpalmer4

OpenStudy (whpalmer4):

one last question :-) \[\log a- \log b = \log(\frac{a}{b})\]

OpenStudy (anonymous):

you dnt have to help @whpalmer4 That was already pre typed

OpenStudy (anonymous):

A. In x^3/c^2 B. ln(x^3 + c^2) C. ln(x^3 - c^2) D. lnx^3c^2

OpenStudy (john_es):

The first option is correct.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you agree @whpalmer4 ?

OpenStudy (whpalmer4):

well, \[\log x^3 = 3\log x\]and \[\log c^2 = 2\log c\]so \[3\log x - 2\log c = \log x^3-\log c^2\]and subtracting logs is like taking the log of a fraction, so \[\log x^3-\log c^2 = \log(\frac{x^3}{c^2})\] So yes, I agree.

OpenStudy (anonymous):

ok great ^_^